# Large Language Models as Urban Residents: An LLM Agent Framework for Personal Mobility Generation

### v1

@@ 5.1 Experimental Setup
Retail employee: typically travel to and from educational institutions at similar times.
Metrics. To evaluate the generation performance, various features related to individual activity are used as metrics: (1) Step Distance (SD)Yuan et al. (2022): This metric evaluates the spatial pattern of an individual’s activities by measuring the distance between two consecutive locations in a trajectory. (2) Step Interval (SI)Yuan et al. (2022): This metric evaluates the temporal pattern of an individual’s activities by measuring the time interval between two successive locations on an individual’s trajectory. (3) ST-ACT: This metric captures and examines the spatial-temporal distribution of activities for each trajectory, focusing on the patterns of individual activities characterized by activity type and timing (e.g., activity type, time). It provides insight into how activities are distributed over space and time. (4) ST-LOC: This metric provides a granular perspective on the generated activities by assessing the spatial-temporal distribution of visited locations within each trajectory, including geographical coordinates and timestamps. It enables a detailed analysis of where and when activities occur.
After extracting the above features from both the generated and real activity trajectory data, Jensen-Shannon divergence (JSD) is employed to quantify the discrepancy between the generated and the actual ones. Lower JSD is preferred.
Baselines. The following models are implemented to serve as baselines: (1) Markov-based mechanic model (MM) Pappalardo and Simini (2018), (2) LSTM-based prediction model (LSTM), (3) Attentional Recurrent Network-based prediction model (DeepMove) Feng et al. (2018).
[…]
5.2 Experimental Results and Analysis
Generative Performance Validation (RQ1, RQ2). The performance evaluation involves analyzing generation results in three distinct settings: (1) Generating normal trajectories based on normal historical trajectories, specifically using data from 2019, a period unaffected by the pandemic. (2) Generating abnormal trajectories based on abnormal historical trajectories, utilizing data from 2020, a year marked by the pandemic. (3) Generating abnormal trajectories based on normal historical trajectories, using data of 2019 and 2021. Through these settings, the efficiency of aligning real-world data, along with the effectiveness of identifying patterns and motivations, can be validated. Furthermore, we expect to examine the models’ generalization capabilities across varying data situations. The results of these evaluations are detailed in the metrics reported in Tables 3 – 5. Through the comparison, it can be observed that although LLMob may not excel in replicating spatial features precisely (such as SD), it demonstrates superior performance in handling temporal aspects. Notably, when considering spatial-temporal features (like ST-ACT and ST-LOC), LLMob’s performance is markedly better. Specifically, we suggest that the pronounced advantage in terms of ST-ACT can be attributed to the LLM agents’ tendency to accurately replicate the motivation behind individual activity behaviors. For instance, they may recognize patterns like a person’s habits to have breakfast in the morning, without being restricted to a specific restaurant. This phenomenon highlights the enhanced semantic understanding capabilities of the LLM agents.
Table 3: Performance (JSD) of normal trajectory generation based on normal historical data.
Models SD SI ST-ACT ST-LOC
MM 0.018 0.276 0.644 0.681
LSTM 0.017 0.271 0.585 0.652
DeepMove 0.008 0.153 0.534 0.623
LLMob-E 0.053 0.046 0.125 0.559
LLMob-L 0.049 0.054 0.136 0.570
LLMob-L w/o $\mathcal{P}$ 0.061 0.080 0.270 0.600
LLMob-E w/o $\mathcal{P}$ 0.055 0.069 0.223 0.530
[…]
Table 4: Performance (JSD) of abnormal trajectory generation based on abnormal historical data.
Models SD SI ST-ACT ST-LOC
MM 0.041 0.300 0.629 0.682
LSTM 0.016 0.286 0.563 0.655
DeepMove 0.011 0.173 0.548 0.668
LLMob-E 0.056 0.043 0.127 0.615
LLMob-L 0.057 0.051 0.124 0.609
LLMob-L w/o $\mathcal{P}$ 0.072 0.081 0.286 0.641
LLMob-E w/o $\mathcal{P}$ 0.059 0.081 0.252 0.573
[…]
Table 5: Performance (JSD) of abnormal trajectory generation based on normal historical data.
Models SD SI ST-ACT ST-LOC
MM 0.039 0.307 0.644 0.681
LSTM 0.035 0.282 0.585 0.653
DeepMove 0.013 0.173 0.534 0.623
LLMob-E 0.072 0.056 0.117 0.536
LLMob-L 0.064 0.051 0.124 0.531
LLMob-L w/o $\mathcal{P}$ 0.073 0.091 0.248 0.580
LLMob-E w/o $\mathcal{P}$ 0.065 0.079 0.209 0.561
[…]
Figure 4: Daily activity frequencies comparison.
Figure 5: Evolving of accumulative counts, Arts & Entertainment.
Figure 6: Evolving of accumulative counts, Professional & Other Places.
Explore Real-World Application (RQ3). We are interested in how LLMob can elevate the social benefits, particularly in the context of urban mobility. To this end, we propose an example of leveraging the flexibility and intelligence of the LLM in understanding semantic information. In particular, we enhance the original setup by incorporating an additional prompt to provide a context for the LLM agent, enabling it to plan activities during specific circumstances. For example, a “pandemic” prompt is as follows: Now it is the pandemic period. The government has asked residents to postpone travel and events and to telecommute as much as possible.
By integrating this prompt, we introduce an external factor – the pandemic situation and the government’s response – into LLMob. In this way, we can observe and analyze the impact of external elements, such as the ongoing pandemic and the government’s measures, on urban mobility and related social dynamics. We use the activity trajectory data during the pandemic (e.g., 2021) as ground truth. As shown in Fig. 4, “LLMob-L” serves as the activity trajectory generator, and the one augmented with the pandemic prompt is denoted as LLMob-L with “Pandemic” Prompt. The influence of the external factor on the simulation is clear: there is a significant decrease in activity frequency when the prompt, which semantically discourages activities prone to spreading the disease, is applied. Additionally, from a spatial-temporal perspective, from spatial-temporal perspective, two major activities (e.g., Arts & Entertainment and Professional & Other Places) are selected to observe the behavior, as shown in Figs. 5 and  6. These activities are particularly insightful as they encapsulate the impact of the pandemic on the work-life balance and daily routines of residents.
Specifically, “LLMob-L with Pandemic Prompt” reproduces a more realistic spatial-temporal activity pattern. This enhanced realism in the generation is attributed to the integration of prior knowledge about the pandemic’s effects and governmental responses, allowing the LLM agent to behave in a manner that aligns with actual behavioral adaptations. For instance, the reduction in Arts & Entertainment activities reflects the closure of venues and social distancing guidelines, while changes in Professional & Other Places activities indicate shifts toward remote work and the transformation of professional environments. Intuitively, prompting LLM agents to generate activities based on various priors shows great potential in real-world applications. The utility of such a conditioned generative approach, coupled with the reliable generated results, can significantly alleviate the workload of urban managers. We suggest that this kind of workflow can simplify the analysis of urban dynamics and aid in assessing the potential impact of urban policies.
6 Conclusion
Contribution. This study is believed to be the first urban mobility simulation empowered by LLM on real-world data. Our innovative framework leverages activity patterns and motivations to direct LLM agents in emulating urban residents, facilitating the generation of interpretable and efficient individual activity trajectory. Extensive experimental studies based on real-world data are conducted to validate the proposed framework. Further exploration demonstrated the promising capabilities of LLM to improve urban mobility management.
Social Impact. Leveraging artificial intelligence to enhance societal benefits is increasingly promising, especially with the advent of high-capacity models such as LLMs. This study explores one of the potential avenues for applications using LLMs as a reliable agent to simulate specific scenarios to assess the effects of external factors, such as government policies. The introduced framework offers a flexible approach to enhance the reliability and efficiency of LLMs in simulating urban mobility.
Acknowledgements
This work is supported by JSPS Kakenhi 22H03903, 23H03406, 23K17456, and JST CREST JPMJCR22M2 and SPRING JPMJSP2108.
Ethical Statement
Despite the use of GPT for polishing the paper, we are responsible for all content. The data is anoynimized. To the best of our knowledge, we find no ethical issues.
Appendix
Appendix A Prompts
where $<$INPUT 0$>$ and $<$INPUT 1$>$ are replaced by the candidate personas. $<$INPUT 2$>$ and $<$INPUT 3$>$ will be replaced by the activity habits extracted from historical data. Specifically, we formulate the $<$INPUT 2$>$ in the following format:
$<$INPUT 2$>$ format
During $<$weekend or weekday$>$, you usually travel over $<$distance$>$ kilometers a day, you usually begin your daily trip at $<$time of first daily activity$>$ and end your daily trip at $<$time of last daily activity$>$. you usually visit $<$location of first daily activity$>$ at the beginning of the day and go to $<$location of last daily activity$>$ before returning home.
where $<$INPUT 0$>$ is the candidate pattern. $<$INPUT 1$>$ is the daily activities plan for evaluation.
where $<$INPUT 0$>$ is replaced by the selected pattern. $<$INPUT 1$>$ is replaced by the number of days from the date to plan activities. $<$INPUT 2$>$ is the historical activities corresponding to the chosen date.
where $<$INPUT 0$>$ is replaced by the selected pattern. $<$INPUT 1$>$ Is replaced by the retrieved historical activities.
[…]
where $<$INPUT 0$>$ is replaced by the selected pattern. $<$INPUT 1$>$ Is replaced by the retrieved motivation. $<$INPUT 2$>$ Is replaced by the most frequently visited locations.
Pandemic
Now it is the pandemic period. The government has asked residents to postpone travel and events and to telecommute as much as possible.
Appendix B Learning-Based Motivation Retrieval
Table 6: Architecture of the score approximator.
Layer Input Size Output Size Notes
Input Layer 3 64 Linear
Activation - - ReLU
Output Layer 64 1 Linear
We include the day of the year for the query date, whether it shares the same weekday as the reference date, and whether both the query and reference dates fall within the same month as input features. Settings for the learning process are as follows: Adam is used as the optimizer, batch size is 64, learning rate is 0.002, and the number of negative samples is 2.

### v2

@@ C.4 Personas
Retail employee: travel patterns might include shifts that
start late in the morning and end in the evening.
Appendix D Additional Experimental Results
D.1 Examples of Identified Patterns
The patterns are extracted and identified during the first phase in our framework. We report some examples of the identified patterns in our experiments as follows, which correspond to the 10 personas in Table 4.
Pattern of a student You are a student in this urban neighborhood. You typically travel to and from educational institutions at similar times.Your daily routine as a student in an urban neighborhood revolves around traveling to and from educational institutions. On weekdays, you cover a distance of over 10 kilometers, starting your daily trip at 12:00 and concluding it at 20:20. At the beginning of the day, you usually visit Park#2457 and then head to Grocery Store#648 before returning home. Weekends follow a similar pattern, with you traveling over 10 kilometers a day, but starting your daily trip at 13:40 and ending it at 19:20. Like on weekdays, you begin by visiting Park#2457 and then proceed to Grocery Store#648 before heading back home. Throughout the week, you have specific places You visit at fixed times, such as Grocery Store#648 at 20:00, Coffee Shop#571 at 09:30, Park#2457 at 08:30, Irish Pub#21 at 20:00, and Bookstore#313 at 16:00. Additionally, there are other locations You occasionally visit, including Chinese Restaurant#168, Supermarket#802, Park#4719, Tea Room#530, Bookstore#336, Pastry Shop#240, Park#2898, Discount Store#807, and Electronics Store#530. By visiting educational institutions at similar times, you ensure punctuality and consistency in your studies. Additionally, your visits to the park and various stores serve as a way to relax, replenish supplies, and indulge in leisure activities. Overall, your daily routine is structured to fulfill both your educational and personal requirements efficiently.
Pattern of a teacher You are a teacher in this urban neighborhood. You typically travel to and from educational institutions at similar times.Your daily routine as a teacher in an urban neighborhood revolves around your regular travel to and from educational institutions. On weekdays, you cover a distance of over 60 kilometers a day, starting your journey at 11:50 in the morning and concluding it at 17:50 in the evening. At the beginning of the day, you usually visit Rest Area#1533 before heading to the Housing Development#101, where you spend a significant amount of time before returning home. During weekends, your travel distance reduces to 50 kilometers a day, with your daily trip commencing at 11:20 and ending at 18:00. On weekends, you follow a different routine, starting your day by visiting Shopping Mall#1262 and then proceeding to Motorcycle Shop#149 before heading back home. Additionally, there are certain locations You occasionally visit, such as Convention Center#101, Food Court#559, Motorcycle Shop#354, and Sports Bar#56, among others. Your daily routine is motivated by the need to fulfill your teaching responsibilities and ensure that you are present at the educational institutions You serve. The specific locations You visit, whether it’s the Rest Area#1533 or Shopping Mall#1262, play a role in providing you with necessary resources, relaxation, or opportunities to engage in personal interests. Overall, your routine is structured to maintain a balance between professional commitments and personal needs.r relevant places, you can effectively serve the needs of the urban neighborhood and contribute to its smooth functioning.
Pattern of a office worker You are a office worker in this urban neighborhood. You have a fixed morning and evening commute, often heading to office districts or commercial centers.Your daily routine as an office worker in an urban neighborhood is quite structured. On weekdays, you travel over 30 kilometers a day, starting your daily trip at 06:50 and ending it at 20:20. Your routine usually begins with a visit to Platform#212, followed by heading to Soba Restaurant#955 before returning home. You also have certain places You visit at specific times, such as Hospital#1255 at 10:30, Platform#212 at 06:30, Soba Restaurant#955 at 22:30, Public Bathroom#76 at 08:30, and Platform#1068 at 05:00. During weekends, your daily travel distance increases to over 40 kilometers, and you start your trip at 07:00, ending it at 20:10. However, the pattern remains the same, starting with a visit to Platform#212 and then going to Soba Restaurant#955 before returning home.
Pattern of a visitor You are a visitor in this urban neighborhood. You tend to travel throughout the day, often visit attractions, dining areas, and shopping districts.Your daily routine as a visitor in this urban neighborhood involves traveling extensively and exploring various attractions, dining areas, and shopping districts. On weekdays, you cover a distance of over 40 kilometers, starting your day at 11:00 and concluding it at 17:10. You have a consistent pattern of visiting Ramen Restaurant#4841 at the beginning of the day and then heading to Town#373 before returning home. Weekends are slightly different, with a shorter distance of around 30 kilometers covered. You begin your day at 14:00 and end it at 16:50. During weekends, you start by visiting Ramen Restaurant#3773 and then proceed to Town#373 before heading back. There are certain locations that you frequently visit, such as Arcade#929 at 13:30, Bowling Alley#306 at 11:00, Arcade#408 at 12:00, and Grocery Store#2094 at 12:30. Additionally, you sometimes visit other places like Electronics Store#562, Comic Shop#5, Video Game Store#8, and many others. Your routine is motivated by your desire to explore and experience the various attractions, cuisines, and shopping opportunities available in this urban neighborhood. You find joy in discovering new places, trying different foods, and immersing yourself in the vibrant atmosphere of this bustling area.
Pattern of a night shift worker You are a night shift worker in this urban neighborhood. You travel to work in the evening and return home in the early morning.Your daily routine as a night shift worker in an urban neighborhood revolves around traveling a considerable distance. On weekdays, you cover over 70 kilometers a day, starting your daily trip at 09:20 and concluding it at 22:30. The routine begins with a visit to Toll Booth#194, followed by a stop at Supermarket#3823 before heading home. During weekends, the pattern remains the same, with the only difference being that you commence your journey at 10:30 and finish at 22:20. On these days, you visit Toll Booth#812 first and then head to Cafe#962 before returning home. In addition to these regular stops, there are occasional visits to various locations such as Record Shop#81, Lake#600, and Factory#495, among others. The motivation behind this routine is to ensure that you are well-prepared for your night shift, starting the day by taking care of essential tasks like stopping at toll booths and getting groceries. These routines help you maintain a sense of order and efficiency in your daily life, ensuring that you are ready for work and able to relax and enjoy your free time.
Pattern of a remote worker You are a remote worker in this urban neighborhood. You may have non-standard travel patterns, often visit coworking spaces or cafes at various times.Your daily routine as a remote worker in an urban neighborhood involves traveling a considerable distance, both during weekdays and weekends. On weekdays, you typically travel over 40 kilometers a day, starting your daily trip at 14:00 and ending it at 19:20. Your routine usually begins with a visit to Convention Center#101, followed by a stop at Supermarket#1593 before returning home. During weekends, your travel distance is slightly less, around 20 kilometers a day, and you begin your daily trip at 11:50, concluding it at 16:30. On weekends, you usually start by visiting Discount Store#884 and then head to Supermarket#1689 before returning home. Additionally, you have specific places You visit at certain times, including Supermarket#1593 at 17:00, Hobby Shop#516 at 13:30, Shopping Mall#1262 at 15:00, Hobby Shop#168 at 14:00, and Exhibit#461 at 11:30. Occasionally, you also visit other locations such as Shopping Mall#1073, Bookstore#14, Shrine#2783, and many more. Your motivation for this routine is to have a flexible work environment, utilizing coworking spaces and cafes, while also fulfilling your daily needs and exploring different places within the urban neighborhood.
Pattern of a service industry worker You are a service industry worker in this urban neighborhood. You might travel outside of standard business hours, often later in the evening or at night.Your daily routine as a service industry worker in an urban neighborhood is quite busy and revolves around your work schedule. On weekdays, you travel over 10 kilometers a day, starting your daily trip at 07:20 and ending it at 20:40. The day usually begins with a visit to Historic Site#2176, followed by a stop at Convenience Store#3385 before returning home. During weekends, your travel distance decreases to 0 kilometers a day, starting your daily trip at 09:20 and ending it at 20:10. Similar to weekdays, you start your day by visiting Historic Site#2176 and then go to Public Art#99 before returning home. Additionally, there are certain locations You sometimes visit, including various convenience stores, restaurants, shopping malls, and other establishments. Your motivation for this routine is primarily driven by your work commitments and the need to fulfill your responsibilities as a service industry worker. It is essential for you to visit specific places, such as convenience stores and historic sites, to ensure You have the necessary supplies and maintain a well-rounded understanding of the neighborhood. Overall, this routine enables you to efficiently navigate your urban neighborhood and fulfill your professional obligations.
Pattern of a public service official You are a public service official in this urban neighborhood. You often work in shifts, leading to varied travel times throughout the day and night.Your daily routine as a public service official in an urban neighborhood revolves around your shifts, which result in different travel times. On weekdays, you typically cover over 60 kilometers a day, starting your daily trip at 12:50 and concluding it at 19:00. At the beginning of the day, you make it a point to visit Convenience Store#3042, and before heading home, you stop by Convenience Store#3702. During the weekends, you travel around 50 kilometers daily, commencing your journey at 10:00 and wrapping it up at 18:30. On weekends, your routine involves visiting Platform#511 in the morning and then heading to Platform#670 before returning home. Additionally, there are a few locations You occasionally visit, such as Platform#1135, Home Service#244, and Convenience Store#6014, among others. The motivation behind your daily routine is to ensure that you cover the necessary ground, making essential stops at various locations to fulfill your duties as a public service official. By visiting convenience stores, platforms, home services, and other relevant places, you can effectively serve the needs of the urban neighborhood and contribute to its smooth functioning.
Pattern of a fitness enthusiast You are a fitness enthusiast in this urban neighborhood. You often travel early in the morning, in the evening, or on weekends to fitness centers or parks. Every weekday, you travel over 70 kilometers a day, starting your daily trip at 10:30 and ending it at 19:50. Your routine begins with a visit to Toll Booth#34, where you kickstart your day. After that, you head to Recreation Center#18 to engage in your fitness activities before returning home. On weekends, your daily travel distance is slightly less, around 60 kilometers. You start your trips at 12:40 and end them at 21:00. The weekend routine starts with a visit to Convenience Store#5940, where you grab some essentials for the day. Then, you head to Convenience Store#8965 before finally returning home. Throughout the week, you also occasionally visit other locations such as Tunnel#1307, Event Space#104, Shopping Mall#217 and #399, and various toll booths. Your motivation for this daily routine is your passion for fitness and maintaining a healthy lifestyle. You prioritize visiting fitness centers or parks during your trips to ensure that you have dedicated time for exercise. Additionally, you make sure to visit convenience stores for any necessary supplies and toll booths for smooth travel. Overall, your routine revolves around staying active, exploring different locations, and ensuring a well-rounded fitness experience.
Pattern of a retail employee You are a retail employee in this urban neighborhood. Your travel patterns might include shifts that start late in the morning and end in the evening. Every weekday, you embark on a daily journey that spans over 50 kilometers. Your routine begins at 09:30, and you conclude your travels at 17:00. To kickstart your day, you always make a point to visit Soba Restaurant#2105, relishing in their delicious offerings. Before heading home, you make a stop at Indian Restaurant#885, savoring their delectable cuisine. On weekends, your daily travel distance decreases slightly to 40 kilometers. You commence your excursions at 09:10 and wrap them up at 17:20. Your first stop on weekends is Hot Spring#205, where you indulge in relaxation and rejuvenation. Before returning home, you make a detour to Department Store#399, exploring the vast array of products they offer. Additionally, there are several other locations You occasionally visit, such as pharmacies, history museums, shrines, shopping malls, and more. Your motivation for this routine stems from your desire for variety and exploration. By visiting different establishments and places, you are able to experience diverse cuisines, immerse yourself in local culture, and find moments of serenity amidst the bustling city life.
D.2 Examples of Retrieved Motivations and Corresponding Generated Trajectories
The retrieved motivations inspire the agent to plan daily activities that closely align with its specific needs. Here are some examples of retrieved motivations and the corresponding daily activities generated in our experiments.
Example 1 Motivation: Your motivations are primarily driven by your desire for entertainment and relaxation, as you enjoy going to the bowling alley and donburi restaurant during late hours, visiting the supermarket and rest area to fulfill your daily needs and take breaks.   Daily Activity Trajectory Generated: Activities at 2019-12-27: Print Shop (35.697, 139.404) at 09:10:00, Entertainment Place (35.733, 139.434) at 10:10:00, Donburi Restaurant (35.730, 139.435) at 10:30:00, Platform (35.733, 139.435) at 10:50:00, Convenience Store (35.721, 139.397) at 11:20:00, Platform (35.729, 139.392) at 21:30:00, Donburi Restaurant (35.699, 139.412) at 22:10:00, Convenience Store (35.697, 139.416) at 22:40:00.
Example 2 Motivation: Your motivations revolve around fulfilling basic needs such as food, convenience, relaxation, and personal interests, including a fondness for Japanese cuisine, maintaining a healthy lifestyle, and enjoying social gatherings in the evening.   Daily Activity Trajectory Generated: Activities at 2019-12-25: Pet Store (35.699, 139.404) at 08:10:00, Convenience Store (35.697, 139.416) at 13:10:00, Gym (35.697, 139.412) at 13:20:00, Massage Studio (35.698, 139.413) at 14:50:00, Electronics Store (35.699, 139.414) at 15:50:00, Fast Food Restaurant (35.697, 139.413) at 17:10:00, Platform (35.698, 139.415) at 18:10:00, Convenience Store (35.701, 139.393) at 19:10:00.
Example 3 Motivation: Your motivations in the urban neighborhood are primarily driven by the need for sustenance and convenience, ensuring personal safety and security, fulfilling work responsibilities, and satisfying personal preferences for food.   Daily Activity Trajectory Generated: Activities at 2019-12-30: Park (35.779, 139.491) at 06:40:00, Convenience Store (35.831, 139.519) at 07:10:00, Discount Store (36.356, 139.064) at 10:40:00, Assisted Living (36.338, 139.048) at 12:00:00, Bank (36.361, 139.059) at 12:10:00, Home Service (36.352, 139.050) at 15:30:00, Sushi Restaurant (36.373, 139.061) at 19:40:00, Hot Spring (36.340, 139.052) at 20:30:00.
Example 4 Motivation: Your motivations revolve around fulfilling basic needs such as food, convenience, relaxation, and personal interests, including a fondness for Japanese cuisine, maintaining a healthy lifestyle, and enjoying social gatherings in the evening.   Daily Activity Trajectory Generated: Activities at 2019-12-29: Park (35.779, 139.491) at 08:20:00, Event Space (35.630, 139.790) at 10:40:00, Convention Center (35.629, 139.795) at 12:20:00, Convenience Store (35.647, 139.817) at 13:00:00, Soba Restaurant (35.643, 139.820) at 13:10:00, Sporting Goods Shop (35.695, 139.762) at 14:50:00, Kushikatsu Restaurant (35.778, 139.495) at 16:40:00.
Example 5 Motivation: Your motivations revolve around your love for exploring different cuisines, seeking convenience and relaxation through internet cafes, and fulfilling your personal needs by visiting the grocery store.   Daily Activity Trajectory Generated: Activities at 2019-12-31: Internet Cafe (35.723, 140.091) at 07:10:00, Donburi Restaurant (35.581, 140.132) at 12:40:00, Japanese Restaurant (35.369, 140.306) at 17:50:00, Rest Area (35.556, 140.208) at 18:30:00, Supermarket (35.865, 140.024) at 19:40:00.